# Optimal Transport for Measures with Noisy Tree Metric

### v1

@@ C.2 Other Discussions
Moreover, as discussed in (Sato et al., 2022), the duplication is not the problem for comparing performances in noisy environments. Indeed, in our simulations, we keep the same settings, the only difference is the transport distance. We have no assumptions/requirements that the considered document datasets are clean in our simulations. Therefore, the noise in the considered document datasets are not a problem for our simulations, but provides more diversity for settings in our simulations.
@@
We applied the clustering-based tree metric sampling method in (Le et al., 2019) to obtain a tree metric and use it as the original given tree metric without perturbation. We emphasize that there is no further processing on that tree metric without perturbation.
@@
As in the experimental results in Section 5, illustrated in Figures 2, and 3, the time consumptions of the robust OT are comparable with the counterpart TW (i.e., OT with tree metric ground cost) for measures with noisy tree metric. The robust OT with the global approach is a little slower than that of others. This is a stark contrast to other approaches for max-min robust OT problem in general. The proposed novel uncertainty sets of tree metrics play an important role for scalability of robust OT for measures with tree metric (e.g., comparing to the $1$-dimensional OT-based approach for max-min robust OT (Deshpande et al., 2019) where there is no efficient/fast algorithmic approach for it yet.)
@@
A concurrent work has recently published in ArXiv (Yamada et al., 2023) where Yamada et al. (2023) leverage the min-max robust variant of tree-Wasserstein for simplicial representation learning by employing self-supervised learning approach based on SimCLR. Empirically, Yamada et al. (2023) also illustrated the advantages of their proposed method over standard SimCLR and cosine-based representation learning.

### v2

@@ C.2 Other Discussions
Moreover, as discussed in Sato et al. (2022), the duplication is not the problem for comparing performances in noisy environments. Indeed, in our simulations, we keep the same settings, the only difference is the transport distance. We have no assumptions/requirements that the considered document datasets are clean in our simulations. Therefore, the noise in the considered document datasets are not a problem for our simulations, but provides more diversity for settings in our simulations.
@@
We applied the clustering-based tree metric sampling method in Le et al. (2019) to obtain a tree metric and use it as the original given tree metric without perturbation. We emphasize that there is no further processing on that tree metric without perturbation.
@@
As in the experimental results in Section 5, illustrated in Figures 2, and 3, the time consumptions of the robust OT are comparable with the counterpart TW (i.e., OT with tree metric ground cost) for measures with noisy tree metric. The robust OT with the global approach is a little slower than that of others. This is a stark contrast to other approaches for max-min robust OT problem in general. The proposed novel uncertainty sets of tree metrics play an important role for scalability of robust OT for measures with tree metric (e.g., comparing to the one-dimensional OT-based approach for max-min robust OT (Deshpande et al., 2019) where there is no efficient/fast algorithmic approach for it yet.)
@@
After the submission of the main manuscript, a concurrent work was published in ArXiv (Yamada et al., 2023) where Yamada et al. (2023) leverage the min-max robust variant of tree-Wasserstein for simplicial representation learning by employing self-supervised learning approach based on SimCLR. Empirically, Yamada et al. (2023) also illustrated the advantages of their proposed method over standard SimCLR and cosine-based representation learning.